OpenStudy (anonymous):

Multiply: 2/10 x 20 = n * Express the product in simplest form. A. 4 B. 20 C. 40 D. 1/100

OpenStudy (anonymous):

I got A

OpenStudy (anonymous):

thanks

OpenStudy (sushi121212):

The answer is 4 because you have multiply it like this: \[\frac{ 2 }{ 10 } \times \frac{ 20 }{ 1 }\] Which will cancel out to be: \[\frac{ 2 }{ 1 }\times \frac{ 2 }{ 1 }\] Or 4, so the answer is A

OpenStudy (anonymous):

You're welcome